Problem 7 (10 pts. Bonus) How many years will it take an investment to triple its initial value if the monthly interest rate is 1% compounded continuously? (Round up your answer to the highest integer) c) 110 years d) 20 years
Answer #1

Let P = $1 and A = $3 (i.e. the future value is three times the investment)

then if interest is continuously compounded then


A = P . e^{rt} , where t is the frequency in months.


3 = 1.e^{0.01*t}

3 = e^{0.01*t}

ln(3) = ln(e^{0.01*t})

1.0986 = 0.01*t

t = 109.86 \simeq 110 months

We may convert months to year. then number of years = 110/12 = 9.16 year = 9 years (approx)